OpenStudy (anonymous):

Radical 174 simplified

OpenStudy (kc_kennylau):

174=2*87=2*3*29 \(\therefore\) cannot be simplified.

OpenStudy (lieutenantgeneral):

√(174) [already simplified] √(560) ==> √(16*35) ==> 4√35